If a 650 cm ladder is placed against a building at a certain angle, it just reaches a point on the building that is 520 cm above the ground. If the ladder is moved to reach a point 80 cm higher up, how much closer will the foot of the ladder be to the building?
Answer by greenestamps(13200)



With the ladder reaching 520 cm above the ground, we have a right triangle with hypotenuse and leg in the ratio 650:520 = 5:4, so the triangle is a scale model of a 3-4-5 right triangle. The distance from the foot of the ladder to the base of the building is the other leg of the triangle, which is 3*130 = 390 cm.

When the ladder is moved so that it reaches 520+80 = 600 cm up on the wall, the right triangle now has hypotenuse and leg in the ratio 650:600 = 13:12. So now we have a scale model of a 5-12-13 right triangle; the distance from the foot of the ladder to the base of the building is now 5*50 = 250 cm.

So the foot of the ladder is now 390-250 = 140 cm closer to the building.

ANSWER: 140 cm closer